# Vojta’s abc conjecture for entire curves in toric varieties highly ramified over the boundary

### v1

@@ 3. The counting function and the $k$-truncated counting function are defined by
Let $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map. Let $D$ be a hypersurface in ${\mathbb{P}}^{n}({\mathbb{C}})$. Let ${\mathcal{Y}}_{f}:={\mathbb{C}}\backslash f^{-1}(D)$ with the smooth exhaustion function $\tau:=\log\sigma$ given above. Then
@@
    $0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\log r+O(1),$
    for $r\geq 1$, where $\widetilde{f}$ the restriction of $f$ on $\mathcal{Y}_{f}$.
@@
    For $r>1$,
    $\mathfrak{X}^{+}_{\sigma}(r)\leq N^{(1)}_{f}(r,D)+C\log r,$
    where $C$ is a constant independent of $f$ and $D$.
We outline a proof here.
@@
Let ${\mathcal{E}}:=f^{-1}(D)$ and write :=${\mathcal{E}}:=\{a_{j}\}_{j=1}^{\infty}$. Denote by $B_{r}^{\sigma}$ and $S_{r}^{\sigma}$ the $\sigma$-ball of radius $r$ and its boundary. From (3.2), $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}$ ($t\geq 1$) and
${\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}D(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}~{}~{}~{}~{}~{}~{}~{}~{}(t\geq{3\over 2}).$
@@
$0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\int_{{3\over 2}}^{r}{dt\over t}\int_{\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\cap{\mathbb{D}}_{t}}f^{*}\omega_{FS}+O(1)$
@@
$=\int_{1}^{r}{dt\over t}\left(\sum_{j=1}^{\infty}\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}\right)+O(1),~{}~{}~{}~{}(r\geq 1).$
Recall in (3.1) that the radius $r_{j}>0$ can be chosen arbitrarily small. For our purpose, for each $j\geq 1$, we choose $r_{j}>0$ sufficiently small so that $\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}<2^{-j}$. Hence the above estimate yields
$0\leq T_{f}(r)-T_{\widetilde{f},\mathcal{Y}_{f}}(r)\leq\log r+O(1),~{}~{}~{}~{}(r\geq 1).$
This proves (i).
To prove (ii), using the above computations and following the argument in [16, Proposition 3.3, pp. 32-33] (see also [2, Line 3 on page 12]), the weighted Euler characteristic satisfies
$\mathfrak{X}^{+}_{\sigma}(r)=\mathfrak{X}_{\sigma}(r)=\int_{1}^{r}\#\{j:|a_{j}|<t\}{dt\over t}+O(\log r).$
Note that #$\#\{{\mathbb{D}}_{t}\cap{\mathcal{E}}\}$ is exactly $n_{f}^{(1)}(D,t)$, which is finite. Since $B^{\sigma}_{r}\subset{\mathbb{D}}_{r}$ for $r\geq 1$,
$\mathfrak{X}^{+}_{\sigma}(r)\leq N^{(1)}_{f}(r,D)+C\log r.$
This completes the proof of (ii). ∎
3.2. Nevanlinna theory on parabolic Riemann surfaces
Let ${\mathcal{Y}}$ be a parabolic open Riemann surface. The following is the starting point of the parabolic Nevanlinna theory.
Lemma 9 (Parabolic Logarithmic Derivative Lemma, Theorem 3.7 in [16] and Proposition 5.1 in [10]).
Let $f$ be a meromorphic function on ${\mathcal{Y}}$. For any $\delta>0$, one has
$m_{f^{\prime}/f,{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{f,{\mathcal{Y}}}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r))+O(1).$
In particular, it gives the following
@@ 3.2. Nevanlinna theory on parabolic Riemann surfaces
If $u$ is a nowhere vanishing holomorphic function on $\mathcal{Y}$ ((((i.e. $u$ is a unit)))), then, for any $\delta>0$, one has
$T_{u^{\prime}/u,{\mathcal{Y}}}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{u,{\mathcal{Y}}}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r))+O(1).$
We recall the following theorem which is an extension of H. Cartan’s theorem. It was from Theorem 6.3 in [10], although the statement is slightly different.
Theorem 11 (Theorem 6.3 in [10]).
Let $f=(f_{0},\dots,f_{n}):{\mathcal{Y}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map whose image is not contained in any proper linear subspaces of ${\mathbb{P}}^{n}({\mathbb{C}})$, where $f_{0},\dots,f_{n}$ are holomorphic functions on ${\mathcal{Y}}$ without common zeros. Let $H_{1},\dots,H_{q}$ be hyperplanes on $\mathbb{P}^{n}(\mathbb{C})$. Then, for $\epsilon>0$,
$\int_{S(r)}\max_{J}\sum_{k\in J}$ $\lambda_{H_{k}}(f)d\mu_{r}+N_{W,\mathcal{Y}}(0,r)$
$\leq_{\rm exc}(n+1+\epsilon)T_{f,{\mathcal{Y}}}(r)+\left(\frac{n(n+1)}{2}+2\epsilon\right)(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+O(1),$
where $W$ is the Wronskian of $f$, and the maximum is taken over all subsets $J$ such that $H_{j},j\in J$, are in general position.
The following theorem can be derived from the proof of [18, Theorem 2.1] and Lemma 9.
@@
Let $f_{0},\dots,f_{n}$ be entire functions on $\mathcal{Y}$ without common zeros. Assume that $f_{n+1}$ is the holomorphic function on ${\mathcal{Y}}$ such that $f_{0}+\dots+f_{n}+f_{n+1}=0$. If $\sum_{i\in I}f_{i}\neq 0$ for any proper subset $I\subset\{0,\dots,n+1\}$, then, for any $\epsilon>0$, we have
$(1-\epsilon)T_{f_{j}/f_{i},{\mathcal{Y}}}(r)\ \leq_{\operatorname{exc}}\sum_{i=0}^{n+1}N_{f_{i},{\mathcal{Y}}}^{(n)}(0,r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)$
for any pair $0\leq i,j\leq n$.
3.3. The moving GCD theorem for parabolic Riemann surfaces

### v2

@@ 3.2. The characteristic functions in the parabolic setting
Let $f=(f_{0},\ldots,f_{n}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic curve, where $f_{0},\ldots,f_{n}$ are entire functions without common zero. Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be given with $\sigma$ being associated to the discs $\{{\mathbb{D}}(a_{j},2r_{j})\}$. Then we can choose $r_{j}$ small enough such that, for all $r\geq 1$, the following hold:
@@
    $O(1)\leq T_{f}(r)-T^{A}_{f,\sigma}(r)\leq\log r+O(1).$
@@
    $O(1)\leq T_{f}(r)-T_{f,\sigma}(r)\leq\log r+O(1).$
@@
From (3.5), for $t\geq{3\over 2}$, $B_{t}^{\sigma}\subset{\mathbb{D}}_{t}$ and
${\mathbb{D}}_{t}\backslash\left(\bigcup_{j=1}^{\infty}D(a_{j},{3\over 2}r_{j})\right)\subset B_{t}^{\sigma}.$
@@
$0\leq T_{f}^{A}(r)-T^{A}_{f,\sigma}(r)\leq\int_{{3\over 2}}^{r}{dt\over t}\int_{\left(\bigcup_{j=1}^{\infty}{\mathbb{D}}(a_{j},{3\over 2}r_{j})\right)\cap{\mathbb{D}}_{t}}f^{*}\omega_{FS}+O(1)$
@@
$\leq\int_{1}^{r}{dt\over t}\left(\sum_{j=1}^{\infty}\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}\right)+O(1),~~~~(r\geq 1).$
We choose $r_{j}>0$ sufficiently small so that
(3.16) $\int_{{\mathbb{D}}(a_{j},{3\over 2}r_{j})}f^{*}\omega_{FS}<2^{-j}.$
Hence the above estimate yields
$0\leq T_{f}^{A}(r)-T^{A}_{f,\sigma}(r)\leq\log r+O(1),~~~~(r\geq 1).$
By using the fact that $T_{f}(r)=T_{f}^{A}(r)+O(1)$, this completes the proof of (i).
To prove (ii), by using (3.8) with $\upsilon:=\log\|{\bf f}\|$, we get
$T^{A}_{f,\sigma}(r)=\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}dd^{c}\log\|{\bf f}\|=\int_{S_{r}^{\sigma}}\log\|{\bf f}\|d\mu_{r}-\int_{B_{r}^{\sigma}}\log\|{\bf f}\|dd^{c}\tau=T_{f,\sigma}(r)-\int_{B_{r}^{\sigma}}\log\|{\bf f}\|dd^{c}\tau.$
Thus
$|T^{A}_{f,\sigma}(r)-T_{f,\sigma}(r)|\leq\int_{B_{r}^{\sigma}}\log^{+}\|{\bf f}\|dd^{c}\tau+\int_{B_{r}^{\sigma}}\log^{+}(1/\|{\bf f}\|)dd^{c}\tau.$
Since $f_{0},\ldots,f_{n}$ have no common zero, we can argue similarly to the proof of Lemma 10 to choose $r_{j}$ small enough so that there exists a constant $C>0$ with
$\int_{B_{r}^{\sigma}}\log^{+}\|{\bf f}\|dd^{c}\tau+\int_{B_{r}^{\sigma}}\log^{+}\frac{1}{\|{\bf f}\|}dd^{c}\tau\leq C.$
for any $r>0$. By combining above with (i), we prove (ii). ∎
Again, these radii $r_{j}$ in Proposition 12 depends on the map $f$.
3.3. The parabolic Logarithmic Derivative Lemma and H. Cartan’s theorem
To establish the parabolic Logarithmic Derivative Lemma, we introduce the weighted Euler characteristic function (see [16, Definition 3.4]):
$\mathfrak{X}^{+}_{\sigma}(r):=\int_{S^{\sigma}_{r}}\log^{+}\left|d\sigma\left({\partial\over\partial z}\right)\right|^{2}d\mu_{r}.$
Using the properties that the disks ${\mathbb{D}}(a_{j},2r_{j})$ are disjoint and that $\sum_{j\geq 1}r_{j}<+\infty$, we see from [16, Example (2), pp. 32 and 33] (see also [2, (2.10)]) that
$\mathfrak{X}^{+}_{\sigma}(r)=\mathfrak{X}_{\sigma}(r)+O(\log r)=N_{\mathcal{E}}(r)+O(\log r),\quad r>1,$
where $N_{\mathcal{E}}(r):=\int_{1}^{r}n_{\mathcal{E}}(t){dt\over t}$, and $n_{\mathcal{E}}(t)$ counts the number of points in ${\mathcal{E}}\cap\mathbb{D}_{t}$ for $t\geq 1$. In particular, if ${\mathcal{E}}:=f^{-1}(D)$ for a holomorphic map $f:{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ and a divisor $D$ on ${\mathbb{P}}^{n}({\mathbb{C}})$ with $f({\mathbb{C}})\not\subset\mbox{Supp}D$, it follows (see also [2, (4.2)]) that
(3.17) $\mathfrak{X}^{+}_{\sigma}(r)\leq N_{f}^{(1)}(D,r)+O(\log r),\quad r>1.$
The above estimate is crucial to our paper. In [2], the authors consider the case when $N_{f}(D,r)=o(T_{f}(r))$, and therefore $\mathfrak{X}^{+}_{\sigma}(r)$ is a small function with respect to $T_{f}(r)$. In our case, we treat the situation where $N_{f}^{(1)}(D,r)\leq{1\over\ell}T_{f}(r)$ for a sufficiently large $\ell$. In this setting, $\mathfrak{X}^{+}_{\sigma}(r)$ should be regarded as having the same order of growth as $T_{f}(r)$.
Lemma 13 (Parabolic Logarithmic Derivative Lemma).
Let $g$ be a meromorphic function on $\mathbb{C}$. Let $\mathcal{Y}$$\subset{\mathbb{C}}$ be as above. For any $\delta>0$, one has
  1. (i)
    $m_{g^{\prime}/g,\sigma}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{g}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r)).$
  2. (ii)
    $m_{g^{(l)}/g,\sigma}(r)\leq_{\rm exc}{l\over 2}\left((1+\delta)^{2}\log T_{g}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r)\right).$
Proof.
(i) We follow the the approach in [16, Theorem 3.8, pp. 26-28]. Consider the auxiliary form
$\Omega:={1\over|w|^{2}(1+\log^{2}|w|)}{\sqrt{-1}\over 2\pi}dw\wedge d\bar{w},$
introduced in [16, (15), p. 27], which has finite volume on $\mathbb{C}$}. By the change of variable formula, we have
$\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}g^{*}\Omega=\int_{w\in{\mathbb{C}}}N_{g,\sigma}(w,r)\Omega.$
Since $B_{t}^{\sigma}\subset\mathbb{D}_{t}$ for any $t\geq 1$, and by the classical First Main Theorem,
$N_{g,\sigma}(w,r)\leq N_{g}(w,r)\leq T_{g}(r)+O(1),~~~\mbox{for}~r>1.$
Thus
$\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}g^{*}\Omega=\int_{w\in{\mathbb{C}}}N_{g,\sigma}(w,r)\Omega\leq\int_{\mathbb{C}}T_{g}(r)\Omega+\int_{\mathbb{C}}C\Omega.$
Consequently, we obtain
$\int_{1}^{r}{dt\over t}\int_{B_{t}^{\sigma}}g^{*}\Omega\leq C_{0}T_{g}(r)$
for some constant $C_{0}>0$. The rest of the proof follows from [16, Theorem 3.8, p. 28]. This finishes the proof of (i).
To prove (ii), notice that
${g^{(l)}\over g}={g^{(l)}\over g^{(l-1)}}{g^{(l-1)}\over g^{(l-2)}}\cdots{g^{\prime}\over g}.$
Therefore, by (i), we obtain
$m_{g^{(l)}/g,\sigma}(r)\leq\sum_{j=1}^{l}m_{g^{(j)}/g^{(j-1)},\sigma}(r)\leq_{\rm exc}{1\over 2}\left((1+\delta)^{2}\sum_{j=1}^{l}\log T_{g^{(j)}}(r)+\delta l\log r+l\mathfrak{X}^{+}_{\sigma}(r)\right).$
By the classical First Main Theorem,
$T_{g^{(j)}}(r)=m_{g^{(j)}}(r)+N_{g^{(j)}}(r)\leq m_{g^{(j)}/g^{(j-1)}}(r)+m_{g^{(j-1)}}(r)+2N_{g^{(j-1)}}(r)\leq_{\rm exc}3T_{g^{(j-1)}}(r)+O(1).$
Hence
$T_{g^{(j)}}(r)\leq_{\rm exc}3^{j}T_{g}(r)+O(1).$
Therefore, we get
$m_{g^{(l)}/g,\sigma}(r)\leq_{\rm exc}{l\over 2}\left((1+\delta)^{2}\log T_{g}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r)\right).$
In particular, by the definition, the above lemma gives the following:
@@
If $u$ is an entire function such that it has no zeros on $\mathcal{Y}$ ((i.e. $u$ is a unit on $\mathcal{Y}$)), then, for any $\delta>0$,
$T^{M}_{u^{\prime}/u,\sigma}(r)\leq_{\rm exc}{1\over 2}((1+\delta)^{2}\log T_{u}(r)+\delta\log r+\mathfrak{X}^{+}_{\sigma}(r)).$
Using the parabolic Logarithmic Derivative Lemma and the parabolic Jensen’s formula, we can easily derive the following parabolic version of H. Cartan’s theorem.
@@
Let $\{\mathcal{Y}$$\mathbb{C}$$\backslash\mathcal{E}$ with the exhaustion function $\sigma$ constructed above. Let $f=(f_{0},\dots,f_{n}):{\mathbb{C}}\rightarrow{\mathbb{P}}^{n}({\mathbb{C}})$ be a holomorphic map whose image is not contained in any proper linear subspace of ${\mathbb{P}}^{n}({\mathbb{C}})$, where $f_{0},\dots,f_{n}$ are entire functions without common zeros. Let $H_{1},\dots,H_{q}$ be hyperplanes on $\mathbb{P}^{n}(\mathbb{C})$. Then, for $\delta>0$,
$\int_{S_{r}^{\sigma}}\max_{J}\sum_{j\in J}\lambda_{H_{j}}(f)d\mu_{r}+N_{W,\sigma}(0,r)$
$\leq_{\rm exc}(n+1)T_{f,\sigma}(r)+O(\log T_{f}(r))+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+c_{\frac{1}{W},\sigma}(r),$
where $W:=W(f_{0},\dots,f_{n})$ is the Wronskian of $f$ and the maximum is taken over all subsets $J\subset\{1,2,\dots,q\}$ such that $H_{j},j\in J$, are in general position.
Proof.
We follow the proof in [17, Theorem A.5.1.3]. Let $H_{j}=\{{\bf w}\in{\mathbb{P}}^{n}({\mathbb{C}})~|~<{{\bf w},\bf a}_{j}>=0\}$ for $j=1,\dots,q$. Let $T$ be the set of all injective maps $\mu:\{0,1,\dots,n\}\rightarrow\{1,\dots,q\}$ such that ${\bf a}_{\mu(0)},\dots,{\bf a}_{\mu(n)}$ are linearly independent. Then
(3.18) $\int_{S_{r}^{\sigma}}\max_{J}\sum_{j\in J}\lambda_{H_{j}}(f)d\mu_{r}$ $\leq$ $\int_{S_{r}^{\sigma}}\max_{\mu\in T}\sum_{j=0}^{n}\log~\left({\|{\bf f}\|\|{\bf a}_{\mu(j)}\|\over|<{\bf f},{\bf a}>|}\right)d\mu_{r}$
$\leq$ $\int_{S_{r}^{\sigma}}\max_{\mu\in T}\log^{+}{|W(<{\bf f},{\bf a}_{\mu(0)}>,\dots,<{\bf f},{\bf a}_{\mu(n)}>)|\over\prod_{j=0}^{n}|<{\bf f},{\bf a}_{\mu(j)}>|}d\mu_{r}$
$+$ $(n+1)T_{f,\sigma}(r)+\int_{S_{r}^{\sigma}}\log(1/|W(f_{0},\dots,f_{n})|)d\mu_{r}+O(1).$
By applying the logarithmic derivative lemma (Lemma 13) and the parabolic Jensen formula (Lemma 9), the theorem follows. ∎
The following result can be derived from the proof of [19, Theorem 2.1] and Lemma 13.
Theorem 16.
@@
Let $f_{0},\dots,f_{n},f_{n+1}$ be entire functions without common zeros. Let $\{\mathcal{Y}$$\mathbb{C}$$\backslash\mathcal{E}$ with the exhaustion function $\sigma$ constructed above. Assume that $f_{0}+\dots+f_{n}+f_{n+1}=0.$ If $\sum_{i\in I}f_{i}\neq 0$ for any proper subset $I\subset\{0,\dots,n+1\}$, then we have
$T^{M}_{f_{i}/f_{j},\sigma}(r)\leq_{\rm exc}\sum_{l=0}^{n+1}N_{f_{l},\sigma}^{(n)}(0,r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\log r)+c_{\frac{1}{D_{0}},\sigma}(r)+\sum_{\begin{subarray}{c}l=0\\ l\neq j\end{subarray}}^{n+1}c_{f_{l},\sigma}(r)+O\left(\max_{1\leq l\leq n}\log T_{\frac{f_{l}}{f_{0}}}(r)\right),$
where $D_{0}$ is an entire function appeared in the proof of [19, Theorem 2.1] or [17, Theorem A5.2.6]), which is determined in terms of $f_{0},\dots,f_{n}$ only.
Proof.
We first sketch the proof by assuming that $f_{0},\dots,f_{n}$ are linearly independent. By applying Theorem 15 to $f=(f_{0},\dots,f_{n+1})$, we have
$\sum_{j=1}^{n+2}m_{f,\sigma}(H_{j},r)+N_{W,\sigma}(r)\leq(n+1)T_{f,\sigma}(r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+c_{\frac{1}{W},\sigma}(r)+O(\log T_{f}(r)),$
where $H_{i}=\{x_{i-1}=0\},i=1,\dots,n+1$ and $H_{n+2}=\{x_{0}+\cdots+x_{n}=0\}$. By Theorem 11, for $1\leq j\leq n+2$,
$T_{f,\sigma}(r)=m_{f,\sigma}(H_{j},r)+N_{f,\sigma}(H_{j},r)+c_{f_{j-1},\sigma}(r).$
Thus we get,
$T_{f,\sigma}(r)\leq\sum_{j=0}^{n+1}N_{f_{j},\sigma}(0,r)-N_{W,\sigma}(r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+c_{\frac{1}{W},\sigma}(r)+\sum_{j=0}^{n+1}c_{f_{j},\sigma}(r)+O(\log T_{f}(r)).$
It is standard to verify that
$\sum_{j=0}^{n+1}N_{f_{j},\sigma}(0,r)-N_{W,\sigma}(r)\leq\sum_{j=0}^{n+1}N^{(n)}_{f_{j},\sigma}(0,r).$
The proof can be completed by applying the following estimate, valid for any for any $0\leq i\leq n$ and $0\leq j\leq n+1$:
$T_{f,\sigma}(r)$ $=\int_{S_{r}^{\sigma}}\log\max_{0\leq l\leq n+1}|f_{l}(z)|d\mu_{r}=\int_{S_{r}^{\sigma}}\log\max_{0\leq l\leq n+1}(|f_{l}(z)|/|f_{j}(z)|)d\mu_{r}+\int_{S_{r}^{\sigma}}\log|f_{j}(z)|d\mu_{r}$
$\geq\int_{S_{r}^{\sigma}}\log^{+}\left|\frac{f_{i}(z)}{f_{j}(z)}\right|d\mu_{r}+N_{f_{j},\sigma}(0,r)+c_{f_{j},\sigma}(r)$
(3.19) $\geq m_{\frac{f_{i}}{f_{j}},\sigma}(r)+N_{\frac{f_{i}}{f_{j}},\sigma}(\infty,r)+c_{f_{j},\sigma}(r)=T^{M}_{f_{i}/f_{j},\sigma}(r)+c_{f_{j},\sigma}(r).$
Thus
$T^{M}_{f_{i}/f_{j},\sigma}(r)\leq\sum_{j=0}^{n+1}N^{(n)}_{f_{j},\sigma}(0,r)+\frac{n(n+1)}{2}(\mathfrak{X}^{+}_{\sigma}(r)+\delta\log r)+c_{\frac{1}{W},\sigma}(r)+\sum_{\begin{subarray}{c}l=0\\ l\neq j\end{subarray}}^{n+1}c_{f_{l},\sigma}(r)+O\left(\max_{1\leq l\leq n}\log T_{\frac{f_{l}}{f_{0}}}(r)\right).$
Passing from the condition that $f_{0},\dots,f_{n}$ are linearly independent to the situation in which no proper subsum of $f_{0}+\dots+f_{n+1}$ vanishes involves only algebraic arguments. In particular, the resulting statements are obtained from the present ones by changing $W$ to $D_{0}$, the modified Wronskian that appears in the proof of [19, Theorem 2.1]. ∎
3.4. The moving GCD theorem in the parabolic setting